find the equation of a parabola with a vertex (2,-12) that passes through the point (0,0)
Answer by stanbon(75887)

find the equation of a parabola with a vertex (2,-12) that passes through the point (0,0)
---------------------
Plotting the vertex and the point (2,-12), the parabola
must open down or to the right.
----
y = a(x-h)^2 + k with a < 0
Using (2,-12) and (0,0)
----------------------------
-12 = a(2-0)^2 + 0
4a = -12
a = -3
-----
Equation:
y = -3(x-2)^2
===================
Cheers,
Stan H.